Q 3 Altawfik Corporation is estimating the cost function for total cost of production of product A using the two points method. The data collected for the past year is as following:

                                       Number of units        Total

                Quarter            produced                  Costs

                      1                     4,000                $   1,000

                      2                     5,400                     1,280

                      3                     7,000                     1,600

                      4                     9,000                     2,000

            Using the two points methods, choose two points with which you estimate the cost function.
Answer #1

The two-point method is to find out the fixed cost and variable cost for the total cost of production.

The high-low method is used to calculate the variable and fixed cost of a product or entity with mixed costs. It takes two factors into consideration. It considers the total dollars of the mixed costs at the highest volume of activity and the total dollars of the mixed cost at the lowest volume of activity. The total amount of fixed costs is assumed to be the same at both points of activity. The change in the total costs is thus the variable cost rate times the change in the number of units of activity.

KEY TAKEAWAYS

  • The high-low method is a simple way to segregate costs with minimal information.
  • The simplicity of the approach assumes the variable and fixed costs as constant, which doesn't replicate reality.
  • Other cost-estimating methods, such as least-squares regression, might provide better results  1. Identify the high and low activity levels from the data set. 2. Calculate the variable cost per unit   3. Calculate the total fixed cost

As per the question the highest number of units produced in 4th quarter i.e 9000 units and lowest production in the 1st quarter i.e. 4000 units. The total cost in the 1st quarter is $1000 and the 4th quarter is 2000 units.

We use high and low technique:

                                              Cost of high activity level-Cost of lower activity level

Variable cost per unit=        ---------------------------------------------------------------------

                                                Highest units- Lowest units

High technique

VC=9000-4000/2000-1000

     =5000/1000= 5 per unit

Total cost=Fixed cost+ variable cost

9000 = FC+2000x5

Fixed cost = Total cost-variable cost

9000-10000= (1000) negative balance it will not justify for cost calculation.

Low technique

4000-1000x5=1000 negative it will also not justify for calculation.